OpenStudy (anonymous):

Solve for x ln 5 + ln (x-4) = ln x

OpenStudy (anonymous):

\[\begin{align*}\ln5+\ln(x-4)&=\ln x\\ \ln5&=\ln x-\ln (x-4)\\ \ln5&=\ln\frac{x}{x-4}\\ 5&=\frac{x}{x-4} \end{align*}\]